Boating: Francois rode his jet ski 36 miles downstream and then 36 miles back. The round trip took 5 and one fourth hours. Find the speed of the jet ski in still water and the speed of the current if the speed of the current is one seventh the speed of the jet ski in still water?
*I had to write the fractions out because I don't know how to do make improper fractions on the computer.

Answer by lwsshak3(11628)

Boating: Francois rode his jet ski 36 miles downstream and then 36 miles back. The round trip took 5 and one fourth hours. Find the speed of the jet ski in still water and the speed of the current if the speed of the current is one seventh the speed of the jet ski in still water?
..
let x=speed of jet ski in still water
speed of current x/7
Speed downstream=x+current = x+x/7=(7x+x)/7=8x/7
Speed upstream=x-current = x-x/7=6x/7
Travel time = distance/speed
36/(8x/7)+36/(6x/7)=5.25
252/8x+252/6x=5.25
LCD=24x
756+1008=126x
x=1764/126=14
ans:
speed of jet ski in still water=14 mph
speed of current = 1/7 of 14=2 mph